OpenStudy (anonymous):

find the area of the triangle whose sides have the given lengths. a=9, b=12, c=15

OpenStudy (anonymous):

Do you know Heron's formula @Chiomatn93 ? Apply that and you will get your answer

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Heron%27s_formula

OpenStudy (anonymous):

\[s= (9+12+15)/2 =18 now by heron's formula Area=\sqrt{18* (18-9) (18-12)(18-15)}\] \ [\sqrt{18*9*6*3} = \sqrt{2916} = 54 therefore your area is 54

OpenStudy (anonymous):

whats the answer and how do u solve?

OpenStudy (anonymous):

Answer is 54

OpenStudy (anonymous):

@Chiomatn93 , Didn't you see the wiki link I gave above ?

OpenStudy (anonymous):

how do u solve it though? like the steps???

OpenStudy (anonymous):

Commo'n the formula is there, the quantities are given. Now can't you multiply and take the square root by yourself

OpenStudy (anonymous):

See the formula is \[Area = \sqrt{s(s-a)(s-b)(s-c)}\] where \[s= {a+b+c \over 2}\] where a , b, c are the length of the sides of the triangle. First find s. then substitute to get your Area ??

OpenStudy (anonymous):

damn calm down...jeez i will do but i just want an example

OpenStudy (anonymous):

see my reply solution is there

OpenStudy (anonymous):

wheres the answer?

OpenStudy (anonymous):

Ok first you would have to take out the s that is semi perimerter formula for s is S= (a +b +c)/2 here the S is 18 Area of triangle is \[A=\sqrt{s*(s-a)*(s-b)*(s-c)}\] So the area would be now \[A=\sqrt{18*(18-9)*(18-12)*(18-15)}\] \[A=\sqrt{18*5*6*3}\] \[A=\sqrt{2916}\] A=54 Therefore the area of triangle is A=54 which is your answer

OpenStudy (anonymous):

I cannot put it more simply

OpenStudy (anonymous):

thank you! i understand it much more now!